# Better Together? A Field Experiment on

### v2

@@ Abstract
Despite algorithms’ potential to improve public service efficiency, lack of evidence and widespread concerns about equity have impeded their deployment. We conduct a large-scale randomized controlled trial ($N=4,681$) where we provide algorithm support to workers allocating Child Protective Services (CPS) investigations in real time. Relative to humans on their own, access to the algorithm reduced maltreatment-related hospitalizations, including a 29 percent reduction in child injury admissions. Algorithm support also improved equity, reducing harm disproportionately for historically disadvantaged groups, while reducing CPS surveillance of Black children. We further show that humans with algorithm support outperformed the algorithm on its own, using counterfactual exercises designed to be applicable in settings where an algorithm-only treatment arm is infeasible. Finally, we provide suggestive evidence using novel LLM-based methods applied to worker discussion notes on how workers reallocated investigations to children at greater likelihood of harm, irrespective of algorithm-predicted risk.
@@ 1 Introduction
Despite the promise of algorithms to improve allocative efficiency (Ludwig et al., 2024), most Americans oppose the use of algorithmic decision aids in criminal justice, hiring, and consumer-facing contexts (Pew Survey, 2018). Experts are similarly divided on the benefits of algorithms for social policy, with leading concerns being embedded bias (Rambachan et al., 2020) and questionable overrides when humans retain discretion over final decisions (Albright, 2023; Angelova et al., 2022). One proposed reason that algorithms have been less helpful than anticipated is that they have not been deployed to help humans leverage their comparative advantage (Ludwig and Mullainathan, 2021). In particular, humans may retain more information than is available to an algorithm and consider payoffs beyond accuracy, such as equity.
@@
This paper provides the first comprehensive field experimental evidence in a social policy setting that human oversight of algorithms can achieve efficiency and equity gains. We conduct a randomized controlled trial involving hundreds of social workers using an algorithm tool in real time in a challenging, high-stakes policy context: Child Protective Services (CPS). Each year, CPS agencies in the United States decide how to respond to four million calls of alleged maltreatment involving a tenth of the U.S. child population (Children’s Bureau 2021), a group acutely at risk of negative later-life outcomes (Doyle 2007, Helénsdotter 2024). Social workers must decide in a matter of minutes which reported families to investigate based on limited information from a caller’s present concern for the child and the family’s prior history with CPS. Agencies do not have the capacity to visit all families, and failure to investigate a child suffering from abuse or neglect can result in costly continued victimization (Currie and Widom 2010) and, in regular instances, death. Misallocation of CPS investigations appears substantial: the majority of first-time confirmed maltreatment victims were previously reported but not investigated. [^3] Authors’ calculations, California investigations from 2014-2019. Consequently, child welfare practitioners have taken an interest in tools that can help improve workers’ allocation of investigations.
@@
We conducted a large-scale field experiment ($N=4,681$) where, for a randomized set of reported families, worker teams that allocate CPS investigations could consult an algorithm decision support tool in addition to their regular information set. The algorithm presented an informative prediction score of children’s foster care removal risk, ranging from 1 to 20, based on past records already accessible to workers. The tool was only available at this decision stage, tool consultation was voluntary, and workers retained full autonomy in deciding which families to investigate. In practice, workers consulted the algorithm in at least 73 percent of decisions.
@@
First, we find that providing algorithm support to human experts improved child wellbeing. Using a unique linkage of statewide hospitalization records to child welfare administrative data, we show that giving workers access to the algorithm reduced children’s future hospitalizations for conditions commonly associated with maltreatment while holding the overall rate of investigation constant. Specifically, relative to children in the control group, the intervention group experienced a 6 percentage point (29 percent) reduction in injury-related hospitalizations and a 0.06 standard deviation reduction in a combined “harm” index constructed from hospital data that included established proxies for abuse and neglect. Access to the algorithm helped workers shield children from the worst (top one percent) tail of the child harm distribution. As over half of children with hospital-confirmed maltreatment belong to this top percentile of the child harm distribution, access to the algorithm appears to have decreased child maltreatment.
@@
Second, access to the algorithmic tool reduced child disparities in health and CPS surveillance. In the control group, Black, Hispanic, female, and low-income children experienced elevated levels of child harm, consistent with prior studies on child mortality (Currie and Schwandt, 2016) and maltreatment prevalence (Putnam-Hornstein et al., 2021). In light of these historic disparities, some U.S. states have forestalled the adoption of algorithm decision aids due to concerns over embedded bias, underscoring a need for evaluation. [^4] https://www.npr.org/2022/06/02/1102661376/oregon-drops-artificial-intelligence-child-abuse-cases We find that decisions made under human-algorithm interaction halved child harm disparities by race, ethnicity, gender, and socioeconomic status. As in national estimates (Kim et al., 2017), Black children in the control group were investigated at much higher rates than other children, even conditional on underlying risk factors. Consistent with a behavioral framework of automatic and deliberative thinking (Kahneman, 2011; Agan et al., 2022), access to algorithmic information reduced disproportionate surveillance: workers with the tool investigated low-risk-score Black children at rates comparable to other low scoring children, without a distinguishable offsetting increase in harm.
@@
We identify as a key mechanism that algorithm support helped workers to parse through complementary information when deciding which families to investigate. Workers allocated investigations more effectively: investigations increased for harmed children, and decreased for children in better health. Triage improved especially for disadvantaged groups. While the algorithm was trained to predict levels of risk, workers may have been considering the treatment effects of an additional investigation. Using proprietary, anonymized text data written during team decision-making discussions, we show with a novel LLM-guided reverse-regression approach that the algorithm changed workers’ discussion patterns. In exploring how discussions changed, we find that the tool appears to have increased the salience of information complementary to algorithm features, such as family structure and changes in circumstances in light of family history. Gains were concentrated among the set of more complex referrals, including those for which workers expressed uncertainty. This suggests that algorithm support may have reallocated worker effort away from extraneous cognitive tasks of looking for family history to instead focusing on the current allegation in light of family history. Attention given to complementary information improved child outcomes without changing the overall level of investigations or other services, and without increasing investigations and services for children with high algorithm risk scores.
@@
Fourth, we develop two counterfactual estimation strategies to (1) bound how well the algorithm could have performed without human oversight and (2) test whether users should have relied marginally more on the algorithm score. The first approach was necessitated by the legal and practical infeasibility of an algorithm-only treatment arm, as in many high-stakes decision environments. The second assessed the policy relevance of imposing a mandatory investigation threshold. On both of these margins, we find that humans with algorithm support likely outperformed counterfactual algorithm-only decisions in terms of both efficiency and equity, despite the algorithm being highly predictive. Worker sensitivity to equity likely played a role in reducing surveillance disparities for disadvantaged groups – particularly Hispanic and low-income children – compared to an algorithm-only decision rule.
@@
This paper’s main contribution is to provide novel experimental evidence that algorithm support can improve efficiency and equity simultaneously without replacing human discretion. Humans and algorithms are complementary in our context; such synergies may not be as “elusive” as feared by Buçinca et al. (2022). Our experimental findings and counterfactual simulations show that humans paired with algorithms outperformed either party on its own. Algorithms and human experts compensate for one another’s shortcomings with respect to equity: human oversight guarded against potential disproportionate surveillance of Hispanic and low-income children under the algorithm, and the algorithm helped reduce high levels of human surveillance of low-predicted-risk Black children. The latter is congruent with prior work that algorithms can indeed shift human reference points (McLaughlin and Spiess, 2024a). In short, pairing an algorithm trained on likely-biased historical data together with potentially-biased human decision-makers can still reduce CPS disparities and improve welfare (Rambachan and Roth, 2020; Gillis et al., 2021). Our findings align with prior quasi-experimental work showing that experts can partly redress algorithmic biases with additional contextual information and by considering additional objectives (De-Arteaga et al., 2020; Obermeyer et al., 2019; Van Donselaar et al., 2010; Hoffman et al., 2018; Gruber et al., 2020; Angelova et al., 2022; Harris and Yellen, 2024).
@@
This study also adds to our collective understanding of when and how human decision-makers benefit from algorithm support. In light of a pattern of suboptimal human overrides (Ye et al., 2022; Agarwal et al., 2023; Mullainathan and Obermeyer, 2022; Albright, 2023), recent work has focused on how to create more elaborate algorithms that circumvent human biases (McLaughlin and Spiess, 2024b). In contrast, our findings highlight that an alternative approach could be to simplify the algorithm, reminiscent of the literature on information overload. Consistent with Vaccaro et al. (2024), in our setting where workers had a built-in comparative advantage and had clarity on the algorithm’s inputs, they were able to integrate the algorithm predictions effectively into their decision-making. They did so by considering other objectives beyond foster care placement risk (Kleinberg et al., 2024), such as equity, legal constraints on subsequent action, and seemingly identifying those who could benefit from an additional investigation, distinguishing estimated treatment effects from predicted risk levels (e.g., Bhatt et al. 2024). An implication of our findings is that algorithms can improve decisions even when workers do not rotely follow algorithm predictions. The tool was designed for workers’ expressed needs in a time- and resource-constrained environment, and its particular efficacy in more complex and ambiguous cases suggests that algorithms may be effective when they reduce extraneous cognitive load. Beyond its design, the algorithm’s deployment in a group-based setting may have also contributed to its effectiveness (Galton, 1907; Patel et al., 2019; Chiang et al., 2023).
@@
We are among the first to design an randomized controlled trial (RCT) to evaluate how professionals use algorithmic information in a high-stakes social policy setting, complementing an important emerging set of human-AI studies. Discontinuity-based designs have identified effects at a specific cutoff margins (Stevenson and Doleac, 2019; Albright, 2023; Cowgill, 2018), whereas our RCT design reveals heterogeneous effects of algorithm support across the distribution of child risk. Our trial takes place in a real-time, consequential setting involving hundreds of professionals: Imai et al. (2023) only evaluate one judge, and medical RCTs that rely on historical case data (e.g., Agarwal et al. 2023), prospective medical RCTs (see Rajpurkar et al., 2022 for a review), and online experiments (Green and Chen, 2019) may not transpose to complex social policy settings. Using methodologies akin to pre-post designs, Goldhaber-Fiebert and Prince (2019) and Rittenhouse et al. (2024) uncover a decrease in racial disparities in CPS interventions under algorithm support, with less clear findings for welfare. We complement a concurrent evaluation of a similarly-styled algorithmic tool in a different child protection agency by Fitzpatrick et al. (2022), who find a small reduction in team discussion time and some change in investigation rates, but due to COVID-related disruptions and a different randomization strategy are statistically underpowered to detect downstream outcomes. Our paper in contrast examines the impacts on efficiency using external data sources on child health, considers consequences for equity, and can speak more generally to human-algorithm complementarity.
@@ 2 Empirical Context
The goal of the child welfare system, also referred to as Child Protective Services (CPS), is to keep children safe from abuse and neglect.555 Technical definitions of abuse and neglect vary slightly by jurisdiction. Abuse refers to direct harm inflicted on a child by a perpetrator, for example physical or sexual abuse, whereas neglect entails failing to protect a child from danger (e.g., drugs, unsafe home environment) or not tending to a child’s medical or developmental needs (See Figure B2 for examples). In this paper, we use the terms “child welfare”, “Child Protective Services”, and “child protection” interchangeably. This section provides a brief description of the child welfare system, accompanied by a description of the machine learning tool developed to aid decision makers.
@@ 2.1 Child Protective Services in the Evaluation County
Appendix Figure A.1 illustrates how CPS operates in our setting. CPS learns about a child who is at risk of maltreatment through reports (“referrals”) to a child protection hotline. Most referrals come from mandated reporters, including school employees, law enforcement, and medical professionals, who are required by law to report any instances of suspected abuse or neglect. A hotline worker records allegations of maltreatment. Most referrals in the study county are sent to a decision team that meets the following weekday morning.
@@ 2.2 Algorithm Tool
The decision aid tool was developed and implemented by Rhema Vaithianathan and colleagues at the Centre for Social Data Analytics (CSDA). It has received favorable and independent support from the scientific community (Goldhaber-Fiebert and Prince, 2019). Similar to previous iterations in Allegheny County (PA) and Douglas County (CO), the tool predicts a child’s likelihood of being removed from their home (placed in foster care) by Child Protective Services within two years of a referral, an outcome that implies high risk of maltreatment. Random forests and XGBoost were considered for the algorithm, but a LASSO regularized logistic regression specification was chosen for runtime purposes. [^6] All data processing and score generation occurred on state servers. This was a deliberate choice by the implementing partner to protect data privacy. The model was trained on a dataset of past child welfare referrals in the trial county and other statewide counties using historical data from approximately 500,000 child-referrals from 2010-2014. The area under the receiver operating characteristic curve (AUC) of the model is 0.76, which indicates good discrimination. [^7] Although the set of variables used to predict removals in the Colorado model was more limited than in Allegheny County due to differences in data availability, the training sample size was much larger. The model in our setting was trained using the full state instead of just one county (larger $n$ but smaller vector of features $X$). As a result, the AUC of the model in our county is slightly better than in Allegheny County, where the AUC was 0.72 (Internal Allegheny DHS report, 2019).
@@
The model uses several hundred features relating to demographics (excluding race, ethnicity, and disability), past child welfare involvement of household members, as well as limited case management data with social services. Of these, most of the algorithm’s predictive power comes from features that encode child and household members’ levels of past CPS involvement. 70% of our sample indeed had prior family history with CPS. The present referral allegation is not an input to the algorithm, which uses only past referral allegations. Text details for incoming referrals are not uploaded to state servers and cannot be used to build the algorithm tool. The choice of which features to include in the model was made in agreement with local implementing partners. Additionally, nearly all information used by the algorithm could be looked up by social workers prior to the trial and during the trial. The algorithm only summarized the information by giving a numerical prediction of removal. As in the two prior deployments of similar tools, predictions were given to workers in ventile (1-20) risk bins, with a score of 1 assigned to children least likely to be removed, and 20 assigned to children at the highest risk of being removed.

### v3

@@ Abstract
Despite algorithms’ potential to improve public services, adoption has been limited by concerns about effectiveness and equity. We conduct a randomized controlled trial ($N=4,681$) providing real-time algorithm support to Child Protective Services (CPS) workers allocating investigations. Algorithm access reduced maltreatment-related hospitalizations, especially among disadvantaged groups, while reducing CPS surveillance of Black children. Notably, child injury admissions decreased by 21 percent. Workers reallocated investigations toward children at greater likelihood of harm, without mechanically following algorithmic predictions. Discussion notes suggest the algorithm shifted worker attention to complementary information. Counterfactual exercises show that human-algorithm complementarity would outperform algorithmic automation in efficiency and equity.
@@ 1 Introduction
Despite the promise of algorithms to improve allocative efficiency (Ludwig et al., 2024), their diffusion into social policy settings has been slow. A majority of the public are not optimistic about the benefits of AI, for example, in criminal justice settings (“How the U.S. Public and AI Experts View Artificial Intelligence” Pew Survey, 2025). Experts are similarly divided on the benefits of algorithms for social policy, with leading concerns being embedded bias (Rambachan et al., 2020) and questionable overrides when humans retain final discretion over decisions (Albright, 2019; Angelova et al., 2022). One proposed reason that algorithms have been less helpful than anticipated is that they have not been deployed to help humans leverage their comparative advantage (Ludwig and Mullainathan, 2021). In particular, humans may retain more information than is available to an algorithm and consider payoffs beyond accuracy, such as equity.
@@
This paper provides the first comprehensive field experimental evidence in a social policy setting that human oversight of algorithms can achieve efficiency and equity gains. The trial involved hundreds of social workers using an algorithm tool in real time in a challenging, high-stakes policy context: Child Protective Services (CPS). Each year, CPS agencies in the United States decide how to respond to four million calls of alleged maltreatment involving a tenth of the U.S. child population (Children’s Bureau 2021), a group acutely at risk of negative later-life outcomes (Doyle 2007, Helénsdotter 2024). Social workers must decide in a matter of minutes which reported families to investigate based on limited information from a caller’s present concern for the child and the family’s prior history with CPS. Agencies do not have the capacity to visit all families, and failure to investigate a child suffering from abuse or neglect can result in costly continued victimization (Currie and Widom 2010) and, in regular instances, death. Misallocation of CPS investigations appears substantial: the majority of first-time confirmed maltreatment victims were previously reported but not investigated. [^3] Authors’ calculations, California investigations from 2014-2019. Consequently, child welfare practitioners have taken an interest in tools that can help improve workers’ allocation of investigations.
@@
We conducted a large-scale field experiment ($N=4,681$) where, for a randomized set of reported families, worker teams that allocate CPS investigations could consult an algorithm decision support tool in addition to their regular information set. The algorithm presented an informative prediction score of children’s foster care removal risk, ranging from 1 to 20, based on past child welfare records already accessible to workers (70 percent had prior history). The tool was only available at this decision stage, tool consultation was voluntary, and workers retained full autonomy in deciding which families to investigate. In practice, workers consulted the algorithm in at least 73 percent of decisions.
@@
First, we find that providing algorithm support improved child wellbeing. Using a unique linkage of statewide hospitalization records to child welfare administrative data, we show that giving workers access to the algorithm reduced children’s future hospitalizations for conditions commonly associated with maltreatment while holding the overall rate of investigation constant. Specifically, relative to children in the control group, the intervention group experienced a 21 percent reduction in injury-related hospitalizations and a 0.05 standard deviation reduction in a combined “harm” index constructed from hospital data that included established proxies for abuse and neglect. Access to the algorithm helped workers shield children from the worst (top one percent) tail of the child harm distribution. As over half of children with hospital-confirmed maltreatment belong to this top percentile of the child harm distribution, access to the algorithm appears to have decreased child maltreatment.
@@
Second, access to the algorithmic tool did not increase child disparities in health and CPS surveillance, but rather it suggestively reduced them. In the control group, Black, Hispanic, female, and low-income children experienced elevated levels of child harm, consistent with prior studies on child mortality (Currie and Schwandt, 2016) and maltreatment prevalence (Putnam-Hornstein et al., 2021). In light of these historic disparities, some U.S. states have forestalled the adoption of algorithm decision aids due to concerns over embedded bias, underscoring a need for evaluation. [^4] https://www.npr.org/2022/06/02/1102661376/oregon-drops-artificial-intelligence-child-abuse-cases In contrast to these concerns, we find that decisions made under human-algorithm interaction halved child harm disparities by race, ethnicity, gender, and socioeconomic status. Although these estimates are not statistical significant at conventional thresholds, they are informative in a Bayesian framework for ethnicity, gender and socioeconomic disparities. Our findings for Black children warrant more caution as they only comprise a modest four percent of our sample. We do however find that Black children in the control group were investigated at much higher rates than other children, in line with national estimates (Kim et al., 2017), even after conditioning on underlying risk factors. Consistent with a behavioral framework of automatic and deliberative thinking (Kahneman, 2011; Agan et al., 2022), access to the algorithm appears to have reduced disproportionate surveillance: workers with the tool investigated low-risk-score Black children at significantly lower rates comparable to other low-scoring children, without a distinguishable offsetting increase in harm.
@@
We identify as a key mechanism that algorithm support helped workers to parse through complementary information when deciding which families to investigate. Workers allocated investigations more effectively: investigations increased for harmed children, and decreased for children in better health. Triage improved especially for disadvantaged groups. Applying cutting-edge sentence-transformer based methods to proprietary, anonymized text data written during team decision-making discussions, we show that the algorithm changed workers’ discussion patterns. In exploring how discussions changed, we find that the tool appears to have increased the salience of information complementary to algorithm features, such as family structure and changes in circumstances in light of family history. Gains were concentrated among the set of more complex referrals, including those for which workers expressed uncertainty. This suggests that algorithm support may have reallocated worker effort away from extraneous cognitive tasks of looking for family history to instead focusing on the current allegation in light of family history.
Child outcomes improved without changing the overall level of investigations or other services, and without increasing investigations and services for children with high algorithm risk scores. One explanation consistent with our findings is that while the algorithm was trained to predict levels of risk based on history and was informative of future child harm, workers also considered the marginal benefit (treatment effect) of an additional investigation. This distinction is consistent with the algorithm’s emphasis on prior family history clueing workers in to consider changes in child and family circumstances, and to consider whether these changes warranted further intervention. More generally, humans considered a broader set of objectives and had more information at their disposal than was available to the algorithm. The multi-pronged evidence for workers’ improved allocation of investigations reinforces a core insight: algorithms can facilitate better decision-making that goes beyond strictly following the algorithm’s risk score.
@@
These findings raise the question of whether forcing workers to investigate high-algorithm-score children would improve outcomes, and motivate a fourth set of results: We develop two counterfactual estimation strategies to (1) bound how well the algorithm could have performed without human oversight and (2) test whether users should have relied marginally more on the algorithm score. An algorithm-only treatment arm was legally and practically infeasible in this context, as in many high-stakes environments, motivating the first counterfactual approach. The second approach assessed the policy relevance of imposing a mandatory investigation threshold. On both of these margins, we find that humans with algorithm support likely outperformed counterfactual algorithm-only decisions in terms of both efficiency and equity, despite the algorithm being highly predictive. Worker sensitivity to equity likely played a role in reducing surveillance disparities for disadvantaged groups – particularly Hispanic and low-income children – compared to an algorithm-only decision rule.
@@
This paper’s main contribution is to provide novel experimental evidence that algorithm support can improve efficiency and possibly even equity without replacing human discretion. Humans and algorithms are complementary in our context; such synergies may not be as “elusive” as feared by Buçinca et al. (2022). Our experimental findings and counterfactual simulations show that humans paired with algorithms outperformed either party on its own. Algorithms and human experts compensate for one another’s shortcomings with respect to equity: human oversight guarded against potential disproportionate surveillance of Hispanic and low-income children under the algorithm, and the algorithm helped reduce high levels of human surveillance of low-predicted-risk Black children. The latter is congruent with prior work that algorithms can indeed shift human reference points (McLaughlin and Spiess, 2024a). In short, pairing an algorithm trained on likely-biased historical data together with potentially-biased human decision-makers can still reduce CPS disparities and improve welfare (Rambachan and Roth, 2020; Gillis et al., 2021). Our findings align with prior quasi-experimental work showing that experts can partly redress algorithmic biases with additional contextual information and by considering additional objectives (De-Arteaga et al., 2020; Obermeyer et al., 2019; Van Donselaar et al., 2010; Hoffman et al., 2018; Gruber et al., 2020; Angelova et al., 2022; Harris and Yellen, 2024).
@@
This study also adds to our collective understanding of when and how human decision-makers benefit from algorithm support. In light of a pattern of suboptimal human overrides (Ye et al., 2022; Agarwal et al., 2023; Mullainathan and Obermeyer, 2022; Albright, 2019), recent work has focused on how to create more elaborate algorithms that circumvent human biases (McLaughlin and Spiess, 2024b). In contrast, our findings highlight that an alternative approach could be to simplify the algorithm, reminiscent of the literature on information overload. Consistent with Vaccaro et al. (2024), in our setting where workers had a built-in comparative advantage and had clarity on the algorithm’s inputs, they were able to integrate the algorithm predictions effectively into their decision-making. They did so by considering other objectives beyond foster care placement risk (Kleinberg et al., 2024), such as equity, legal constraints on subsequent action, and seemingly identifying children who could benefit from an additional investigation, distinguishing estimated treatment effects from predicted risk levels (e.g., Bhatt et al. 2024). An implication of our findings is that algorithms can improve decisions even when workers do not rotely follow algorithm predictions. The tool was designed for workers’ expressed needs in a time- and resource-constrained environment, and its particular efficacy in more complex and ambiguous cases suggests that algorithms may be effective when they reduce extraneous cognitive load. Beyond its design, the algorithm’s deployment in a group-based setting may have also contributed to its effectiveness (Galton, 1907; Patel et al., 2019; Chiang et al., 2023).
@@
We are among the first to design an randomized controlled trial (RCT) to evaluate how professionals use algorithmic information in a high-stakes social policy setting, complementing an important emerging set of human-AI studies. Discontinuity-based designs have identified effects at a specific cutoff margins (Stevenson and Doleac, 2024; Albright, 2023; Cowgill, 2018), whereas our RCT design reveals heterogeneous effects of algorithm support across the distribution of child risk. Our trial takes place in a real-time, consequential setting involving hundreds of professionals: Imai et al. (2023) only evaluate one judge, and medical RCTs that rely on historical case data (e.g., Agarwal et al. 2023), prospective medical RCTs (see Rajpurkar et al., 2022 for a review), and online experiments (Green and Chen, 2019) may not transpose to complex social policy settings. Using methodologies akin to pre-post designs, Goldhaber-Fiebert and Prince (2019) and Rittenhouse et al. (2024) uncover a decrease in racial disparities in CPS interventions under algorithm support, with less clear findings for welfare. We complement a concurrent evaluation of a similarly-styled algorithmic tool in a different child protection agency by Fitzpatrick et al. (2022), who find a small reduction in team discussion time and some change in investigation rates, but due to COVID-related disruptions and a different randomization strategy are statistically underpowered to detect downstream outcomes. Our paper in contrast examines the impacts on efficiency using external data sources on child health, considers consequences for equity, and can speak more generally to human-algorithm complementarity.
@@ 2 Empirical Context
The goal of the child welfare system, also referred to as Child Protective Services (CPS), is to keep children safe from abuse and neglect. [^5] Technical definitions of abuse and neglect vary slightly by jurisdiction. Abuse refers to direct harm inflicted on a child by a perpetrator, for example physical or sexual abuse, whereas neglect entails failing to protect a child from danger (e.g., drugs, unsafe home environment) or not tending to a child’s medical or developmental needs (See Figure B2 for examples). In this paper, we use the terms “child welfare”, “Child Protective Services”, and “child protection” interchangeably. This section provides a brief description of the child welfare system, accompanied by a description of the machine learning tool developed to aid decision makers.
@@ 2.1 Child Protective Services in the Evaluation County
Appendix Figure A.1 illustrates how CPS operates in our setting. CPS learns about a child who is at risk of maltreatment through reports (“referrals”) to a child protection hotline. Three-fourths of referrals in our setting originate from mandated reporters, including school employees, law enforcement, and medical professionals, who are required by law to report any instances of suspected abuse or neglect. A hotline worker records allegations of maltreatment. Most referrals in the study county are sent to a decision team that meets the following weekday morning.
@@ 2.2 Algorithm Tool
The decision aid tool was developed and implemented by Rhema Vaithianathan and colleagues at the Centre for Social Data Analytics (CSDA). It has received favorable and independent support from the scientific community (Goldhaber-Fiebert and Prince, 2019). Similar to previous iterations in Allegheny County (PA) and Douglas County (CO), the tool predicts a child’s likelihood of being removed from their home (placed in foster care) by Child Protective Services within two years of a referral, an outcome that implies high risk of maltreatment. Foster care was selected as the model’s outcome because – despite some discretion over foster care placement – it is one of the more relatively firm indicators of severe concern available internally to CPS agencies. Random forests and XGBoost were considered for the algorithm, but a LASSO regularized logistic regression specification was chosen for runtime purposes. [^6] All data processing and score generation occurred on state servers. This was a deliberate choice by the implementing partner to protect data privacy. The model was trained on a dataset of past child welfare referrals in the trial county and other statewide counties using historical data from approximately 500,000 child-referrals from 2010-2014. The area under the receiver operating characteristic curve (AUC) of the model is 0.76, which indicates good discrimination. [^7] Although the set of variables used to predict removals in the Colorado model was more limited than in Allegheny County due to differences in data availability, the training sample size was much larger. The model in our setting was trained using the full state instead of just one county (larger $n$ but smaller vector of features $X$). As a result, the AUC of the model in our county is slightly better than in Allegheny County, where the AUC was 0.72 (Internal Allegheny DHS report, 2019).
@@
The model uses several hundred features relating to demographics (excluding race, ethnicity, and disability), past child welfare involvement of household members, as well as limited case management data with social services. [^8] Though the model’s full feature set is public, the subset of features selected by the LASSO is not disclosed, in accordance with an agreement between the tool designers and the county. Of these, CSDA reports that most of the algorithm’s predictive power comes from features that encode child and household members’ levels of past CPS involvement. 71 percent of our sample indeed had prior family history with CPS, and the remaining 29 percent still had some variation in other predictors. The present referral allegation is not an input to the algorithm, which relies primarily on past referral allegations. Text details for incoming referrals are not uploaded to state servers and cannot be used to build the algorithm tool. The choice of which features to include in the model was made in agreement with local implementing partners. Additionally, nearly all information used by the algorithm could be looked up by social workers prior to the trial and during the trial. The algorithm only summarized the information by giving a numerical prediction of removal. As in the two prior deployments of similar tools, predictions were given to workers in ventile (1-20) risk bins, with a score of 1 assigned to children least likely to be removed, and 20 assigned to children at the highest risk of being removed.